Given the following information about an open economy with government draw the AD function and find the equilibrium level of income?
C=100+0.7YD, I=200, G =150, t=A0.2, X=100, Z=0.2
1
Expert's answer
2020-11-27T07:34:38-0500

let A=2 than

"Y=C+I+G+(X-Z)"

"Y=100+0.7Yd+200+150+(100-0.2)"

"Y=549.8+0.7Yd"

Yd=Y-T

"t = A0,2=2\times0.2=0.4"

"Y=549.8+0.7(Y-0.4)=549.52+0.7Y"

"Y-0.7Y=549.52"

Y=1871.33


AD=C+I

C = 100 + 0.7YD

I = 200

AD=100 + 0.7YD+200=300+0.7YD

AD=300+0.7YD